2(a-4)-3(a-2)=4(a+1)+4
Answer by rfadrogane(214)

2(a-4)-3(a-2)=4(a+1)+4
Sol'n:
2a-8-3a+6 = 4a+4+4
-a-2 = 4a+8
4a+a = -2-8
5a = -10
a = -2 -----answer